OpenStudy (anonymous):

solve for w:1/2 =-1/3w-1/5

OpenStudy (anonymous):

like this? \[\frac{ 1 }{ 2 } = \frac{ -1 }{ 3w } - \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

no this\[\frac{ 1 }{2 }= -\frac{ 1 }{ 3 }w-\frac{ 1 }{ 5}\]

OpenStudy (anonymous):

Okay! I got w = \[-\frac{ 20 }{ 10 }\] or \[\approx -2.1000\]